You took a quiz in Mathematics but your teacher did not tell you your exact answer was. Instead,he informed you that 20% of your classmates higher than you. What was your score?

If your IQ is 85, what percentage of people have a higher IQ than you? IQ mean is 100 and the standard deviation is 15.
1
Expert's answer
2020-10-19T17:51:38-0400

Let "X=" IQ score: "X\sim N(\mu, \sigma^2)"

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given "\mu=100, \sigma=15"

If 20% of your classmates higher than you , what is your IQ?


"P(X> x)=1-P(X\leq x)=""=1-P(Z\leq\dfrac{x-100}{15})=0.2"

"P(Z\leq\dfrac{x-100}{15})=0.8"

"\dfrac{x-100}{15}\approx0.841621"

"x\approx112.6"

My score was 112.6.


If my IQ is 85, what percentage of people have a higher IQ than I?


"P(X> 85)=1-P(X\leq 85)=""=1-P(Z\leq\dfrac{85-100}{15})=1-P(Z\leq-1)\approx""\approx1-0.158655=0.841345"

If my IQ is 85, approximately 84% of people have a higher IQ than I.